# thanks: Corresponding authorthanks: Corresponding author

### v1

@@ III RESULTS AND DISCUSSION
at the $1\sigma$ confidence level (C.L.). Our estimate agrees with the Planck 2018 result of $H_{0}=67.4\pm 0.5~{}\mathrm{km~{}s^{-1}~{}Mpc^{-1}}$ Aghanim et al. (2020). However, a mild tension persists at the $2.0\sigma$ level in comparison with the local SH0ES measurement of $H_{0}=73.04\pm 1.04~{}\mathrm{km~{}s^{-1}~{}Mpc^{-1}}$ Riess et al. (2022). Other datasets exhibit the same agreement with the Planck result and a similar tension level with the SH0ES measurements. Specifically, DESY5+SDSS+CC datasets show a $2.1\sigma$ discrepancy, while PantheonPlus+DESI DR2+CC and PantheonPlus+SDSS+CC datasets exhibit $1.6\sigma$ and $1.7\sigma$ discrepancies, respectively. Our results show larger uncertainty than other traditional IDLs Camilleri et al. (2025); Aubourg et al. (2015); Macaulay et al. (2019) due to the substitution of the Gaussian $r_{\mathrm{d}}$ prior with the more uncertain CC measurements.
Notably, the DESY5 data yield a lower $H_{0}$ and a higher $M_{\mathrm{B}}$ compared to the PantheonPlus data, which tends to enlarge the $H_{0}$ tension. This trend is consistent with previous findings reported in Refs. Colgáin and Sheikh-Jabbari (2024); Notari et al. (2024); Adame et al. (2025). A shift in $r_{\mathrm{d}}$ is observed between the DESI DR2 and SDSS BAO datasets, across both supernova samples. This behavior is expected within the IDL framework, where the CC data serve to anchor the BAO measurements and infer the $r_{\mathrm{d}}$, rather than relying on a prior $r_{\mathrm{d}}$. DESI DR2 data yield a higher $r_{\mathrm{d}}$ than SDSS data by approximately $2~{}\mathrm{Mpc}$.
[…]
@@
Figure 2: The constraints on $\eta$ and $p_{\mathrm{age}}$ in the PAge model obtained from DESY5+DESI DR2+CC (green), DESY5+SDSS+CC (blue), PantheonPlus+DESI DR2+CC (purple), and PantheonPlus+SDSS+CC (pink). The DESI DR2 and Planck 2018 results are introduced for comparison. Since $\Lambda\mathrm{CDM}$ lacks one parameter compared to the PAge model, it is mapped as a straight line in the parameter space.
@@
Figure 3: The BAO distance for the PAge (blue), $w_{0}w_{a}\mathrm{CDM}$ (green), and $w\mathrm{CDM}$ (orange) models in DESY5 + DESI DR2 + CC. These distances are normalized by the best-fit $\Lambda\mathrm{CDM}$ model (black) in DESY5+DESI+CC. The data of DESI (purple, filled) are drawn with error bars.
Figure 4: Model comparison across four datasets using Bayesian evidence for the PAge model and the $\Lambda\mathrm{CDM}$ model, evaluated according to the Jeffreys’ scale. We take $\Lambda\mathrm{CDM}$ as the baseline, so positive values indicate a preference for the PAge model in the corresponding dataset.
[…]
@@
After examining how the choice of SNe and BAO datasets influences the cosmological parameters, we now focus on the approximation parameters of the PAge model itself, namely $\eta$ and $p_{\mathrm{age}}$. These two parameters play a central role in offering insights into potential late-time modifications to cosmic expansion. The posterior distributions in the $(p_{\mathrm{age}},\eta)$ plane for the four datasets are shown in Fig. 2. For comparison, previous results from $\Lambda\mathrm{CDM}$ Aghanim et al. (2020) and DESI DR2 Abdul Karim et al. (2025) are mapped into the PAge model. The mapping procedure is detailed in Eqs. (4), (6), (7), (8), and (9). Note that the PAge model contains one additional parameter compared to $\Lambda\mathrm{CDM}$, which causes the mapped region of $\Lambda\mathrm{CDM}$ in the $(p_{\mathrm{age}},\eta)$ space to appear as a line. We adopt the DESI DR2 results for $w_{0}w_{a}\mathrm{CDM}$ and $w\mathrm{CDM}$ models Abdul Karim et al. (2025). As shown in Fig. 2, it is evident that our results deviate from the $\Lambda\mathrm{CDM}$ model, showing a lower value of $\eta$ for the DESY5 dataset. In contrast, for the PantheonPlus samples, different BAO combinations remain consistent with $\Lambda\mathrm{CDM}$.
@@
These findings suggest that the DESI DR2 data may not be fully compatible with the $\Lambda\mathrm{CDM}$ model, which lack the flexibility to capture such rapid variations in the expansion history, thereby hinting at a preference for dynamical dark energy. It is important to emphasize that the PAge model is essentially a calibrated Taylor expansion, and therefore independent of any specific cosmological model. As such, our results serve as a model-agnostic indication of the underlying behavior of the universe. The presence of this trend in our findings suggests that any viable theoretical model must be capable of reproducing this trend to be consistent with the data.
To quantitatively evaluate the performance of the PAge model relative to the $\Lambda\mathrm{CDM}$ model, we perform Bayesian inference across different datasets. Specifically, the Bayes factors $\mathrm{ln}\mathcal{B}_{ij}$ are calculated in logarithmic space,
@@
where $P(D\mid\theta,M)$ is the likelihood of the data $D$ given the parameters $\theta$ and model $M$, $P(\theta\mid M)$ is the prior probability of $\theta$ under $M$, and $P(M)$ is the prior of $M$. The extent of model preference is assessed using the Jeffreys’ scale, as described in Ref. Trotta (2008).
@@
  • $2.51\leq|\mathrm{ln}\mathcal{B}_{ij}|<$ 2.5, weak evidence.
  • $52.5\leq|\mathrm{ln}\mathcal{B}_{ij}|<$ 5, moderate evidence.
  • $105\leq|\mathrm{ln}\mathcal{B}_{ij}|<$ 10, strong evidence.
@@
In Fig. 4, we present the final model comparison results, with the PAge model assessed relative to the fiducial $\Lambda\mathrm{CDM}$ across four datasets. Positive values indicate a preference for the PAge model, while negative values indicate a preference for the $\Lambda\mathrm{CDM}$ model. Specifically, we find $\mathrm{ln}\mathcal{B}_{ij}=9.17,~{}4.81,~{}1.01,~{}\text{and}~{}2.71$ for DESY5+DESI DR2+CC, DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC, respectively. It is clear that the PAge model consistently provides a better fit than the standard $\Lambda\mathrm{CDM}$ across all datasets. In particular, for the DESY5+DESI DR2+CC datasets, the Bayes factor is 9.17, indicating nearly decisive evidence in favor of the PAge model over the $\Lambda\mathrm{CDM}$ model.
@@ IV CONCLUSION
The tension between the early-universe and late-universe measurements of $H_{0}$ has become one of the most significant challenges in modern cosmology. Recently, DESI DR1 Adame et al. (2025) and DR2 Abdul Karim et al. (2025) have further sharpened this tension with addition hints of dynamical dark energy in $3.9\sigma$ and $4.2\sigma$ level. In the context of the $H_{0}$ tension, numerous theoretical models have been proposed, although many models face corresponding no-go arguments. Similarly, the hint of dynamical dark energy has been critically examined through the reanalyzes of the DESY5 SNe data Huang et al. (2025b); Efstathiou (2024); Vincenzi et al. (2025); Gialamas et al. (2025) and the DESI BAO measurements Zheng et al. (2024); Mukherjee and Sen (2025); Liu et al. (2024).
[…]
@@
Mapping specific cosmological models into the PAge parameter space $(p_{\mathrm{age}},\eta)$, our results align with the DESI DR2 hints of dynamical dark energy and deviate from $\Lambda\mathrm{CDM}$. In particular, the reconstructed BAO distance of the PAge model reveals a special trend compared to best-fit $\Lambda\mathrm{CDM}$. It shows a slight enhancement in very low redshift, suppression at intermediate redshift, and an upturn at high redshift, which may suggest that the $\Lambda\mathrm{CDM}$ model is insufficient to fully reconcile the observed data. Furthermore, the Bayesian model comparison provides evidence in favour of the PAge model over $\Lambda\mathrm{CDM}$, with Bayes factors of 9.17, 4.81, 1.01 and 2.71 in DESY5+DESI DR2+CC, DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, PantheonPlus+SDSS+CC respectively. It gives nearly decisive evidence in favour of the PAge model over $\Lambda\mathrm{CDM}$ in the DESY5+DESI DR2+CC combination.
@@
In future work, we aim to explore the data in more detail by performing a redshift-binned analysis. Motivated by the connection between the expansion history and the dark-energy equation of state, we will further investigate the occurrence of unphysical phantom behavior Caldwell et al. (2003); Vikman (2005) using the PAge model. It should be noted that the PAge model is essentially a Taylor expansion, which may suffer from singularities as discussed in Ref. Zhang et al. (2017b) when reconstructing the equation-of-state parameter. We are going to calibrate the Taylor expansion with alternatives beyond the early matter dominated universe used in the PAge model, to avoid the singularities. These efforts may provide new insights into the nature of dark energy and further clarify the physical origin of the $H_{0}$ tension.

### v2

@@ III RESULTS AND DISCUSSION
at the $1\sigma$ confidence level. Our estimate agrees with the Planck 2018 result of $H_{0}=67.4\pm 0.5~\mathrm{km~s^{-1}~Mpc^{-1}}$ Aghanim et al. (2020). However, a mild tension persists at the $2.0\sigma$ level in comparison with the local SH0ES measurement of $H_{0}=73.04\pm 1.04~\mathrm{km~s^{-1}~Mpc^{-1}}$ Riess et al. (2022). Other datasets exhibit the same agreement with the Planck result and a similar tension level with the SH0ES measurements. Specifically, DESY5+SDSS+CC datasets show a $2.1\sigma$ discrepancy, while PantheonPlus+DESI DR2+CC and PantheonPlus+SDSS+CC datasets exhibit $1.6\sigma$ and $1.7\sigma$ discrepancies, respectively. Our results show larger uncertainty than other traditional IDLs Camilleri et al. (2025); Aubourg et al. (2015); Macaulay et al. (2019) due to the substitution of the Gaussian $r_{\mathrm{d}}$ prior with the more uncertain CC measurements.
From DESY5+DESI DR2+CC, we obtain $M_{\mathrm{B}}=-19.36\pm 0.07$, consistent with the Gaussian processes estimate $M_{\mathrm{B}}=-19.31^{+0.09}_{-0.11}$ Favale et al. (2023), with $M_{\mathrm{B}}=-19.303$ derived from tip of the red giant branch with full sample in Hubble Space Telescope or James Webb Space Telescope Li et al. (2025f), and $1.43\sigma$ differing from the SH0ES result $M_{\mathrm{B}}=-19.253\pm 0.027$ Riess et al. (2022). For sound horizon, we find $r_{\mathrm{d}}=145.0^{+4.5}_{-5.2}~\mathrm{Mpc}$, in agreement with Planck result $r_{\mathrm{d}}=147.05\pm 0.30~\rm Mpc$ and $r_{\mathrm{d}}=146.0^{+4.2}_{-5.1},~141.9^{+5.6}_{-4.9}~\mathrm{Mpc}$ reported in Refs. Gómez-Valent (2022); Favale et al. (2023). As shown in Fig. 2, despite the constraints on $M_{\mathrm{B}}$ and $r_{\mathrm{d}}$ being broad and remain $\sim 1\sigma$ consistent with Planck and SH0ES, the value of $H_{0}$ differs from SH0ES by $2.0\sigma$. This higher discrepancy stems from $a_{\mathrm{B}}$ offset between DESY5 and PantheonPlus and difference in $H_{0}r_{\mathrm{d}}$ inferred from DESI and SDSS. As reported in Ref. Efstathiou (2025), DESY5 and PantheonPlus exhibit an offset of $\sim 0.04$ magnitude in their overlap. Moreover, Ref. Huang et al. (2025b) finds that $5a_{\mathrm{B}}$ for DESY5 is higher than that for PantheonPlus. From Eq. (15), at fixed $M_{\mathrm{B}}$, a larger value of $5a_{\mathrm{B}}$ implies smaller $H_{0}$, thereby increasing the tension with SH0ES. As shown in the $H_{0}-M_{\mathrm{B}}$ plane of Fig. 2, holding $M_{\mathrm{B}}$ fixed, DESY5 favors a lower value of $H_{0}$, consistent with Refs. Colgáin and Sheikh-Jabbari (2024); Notari et al. (2024); Adame et al. (2025). Note that the value of $H_{0}r_{\mathrm{d}}$ inferred from DESI is higher than those from SDSS Mukherjee and Sen (2024); Adame et al. (2025), which would, for fixed $r_{\mathrm{d}}$, raise the inferred $H_{0}$. Nevertheless, this effect is smaller than the effect of SNe as shown in $H_{0}-M_{\mathrm{B}}$ and $H_{0}-r_{\mathrm{d}}$ planes across four datasets in Fig. 2. Consequently, although $M_{\mathrm{B}}$ is within $1.4\sigma$ of SH0ES, the inferred $H_{0}$ differs from the SH0ES value by $2.0\sigma$.
Figure 2: Constraints on the PAge model from DESY5+DESI DR2+CC (green), DESY5+SDSS+$C$ (blue, dotted), PantheonPlus+DESI DR2+CC (purple), and PantheonPlus+SDSS+CC (brown, dotted). Planck 2018 (pink) and SH0ES (grey) benchmarks are included to illustrate the $H_{0}$ tension both from the direct $H_{0}$ constraints and from those inferred via $M_{\mathrm{B}}$ and $r_{\mathrm{d}}$. The SH0ES-equivalent $r_{\mathrm{d}}$ and Planck-equivalent $M_{\mathrm{B}}$ are obtained by equating $H_{0}r_{\mathrm{d}}$ and $a_{\mathrm{B}}$, respectively. Four datasets are presented in the $M_{\mathrm{B}}$$H_{0}$ plane to illustrate the shift in the intercept of the magnitude-redshift relation vividly.
@@
Figure 3: The constraints on $\eta$ and $p_{\mathrm{age}}$ in the PAge model obtained from DESY5+DESI DR2+CC (green), DESY5+SDSS+$C$ (blue), PantheonPlus+DESI DR2+CC (purple), and PantheonPlus+SDSS+CC (pink). The DESI DR2 and Planck 2018 results are introduced for comparison. Since $\Lambda\mathrm{CDM}$ lacks one parameter compared to the PAge model, it is mapped as a straight line in the parameter space.
@@
Figure 4: The BAO distance for the PAge (blue), $w_{0}w_{a}\mathrm{CDM}$ (green), and $w\mathrm{CDM}$ (orange) models in DESY5 + DESI DR2 + CC. These distances are normalized by the best-fit $\Lambda\mathrm{CDM}$ model (black) in DESY5+DESI+$C$. The data of DESI (purple, filled) are drawn with error bars.
Figure 5: The reconstructed BAO distance for the PAge (green), MAPAge (red), incorporating DESI DR2 data. The shaded bands denote the corresponding $1\sigma$ uncertainties.
Figure 6: Bayesian model comparison of PAge with $\Lambda\rm CDM$ (left panel) and $w_{0}w_{a}\mathrm{CDM}$ model (right panel). Positive values indicate a preference for the PAge model in each dataset. The impact of prior widths for $\eta$ and $p_{\mathrm{age}}$ is assessed by broadening the original prior width $d$ to $d^{*}=\alpha d$, with $\alpha$ denoting the broadening factor.
@@
After examining how the choice of SNe and BAO datasets influences the cosmological parameters, we now focus on the approximation parameters of the PAge model itself, namely $\eta$ and $p_{\mathrm{age}}$. These two parameters play a central role in offering insights into potential late-time modifications to cosmic expansion. The posterior distributions in the $(p_{\mathrm{age}},\eta)$ plane for the four datasets are shown in Fig. 3. For comparison, previous results from $\Lambda\mathrm{CDM}$ Aghanim et al. (2020) and DESI DR2 Abdul Karim et al. (2025) are mapped into the PAge model. The mapping procedure is detailed in Eqs. (6), (11), (12), (13), and (14). Note that the PAge model contains one additional parameter compared to $\Lambda\mathrm{CDM}$, which causes the mapped region of $\Lambda\mathrm{CDM}$ in the $(p_{\mathrm{age}},\eta)$ space to appear as a line. We adopt the DESI DR2 results for $w\mathrm{CDM}$ and $w_{0}w_{a}\mathrm{CDM}$ models Abdul Karim et al. (2025). As shown in Fig. 3, it is evident that our results deviate from the $\Lambda\mathrm{CDM}$ model, showing a lower value of $\eta$ for the DESY5 dataset. In contrast, for the PantheonPlus samples, different BAO combinations remain consistent with $\Lambda\mathrm{CDM}$.
@@
These findings suggest that the DESI DR2 data may not be fully compatible with the $\Lambda\mathrm{CDM}$ model, which lack the flexibility to capture such rapid variations in the expansion history, thereby hinting at a preference for dynamical dark energy. PAge provides a model-independent description of the underlying behaviors of the late-time Universe. However, it should be noted that the PAge model rests on specific assumptions. Thus, PAge may not capture inhomogeneous cosmologies, nor scenarios that violate the assumptions (i) and (ii).
Then, we fit the MAPAge model to DESY5+DESI DR2+CC instead of PAge, to assess the impact of the chosen truncated order in Eq. (1). Our final constraints are: $\eta_{1}=-0.01\pm 0.09,~\eta_{2}=0.65\pm 0.22,~p_{\mathrm{age}}=0.93\pm 0.01,~H_{0}=67.57\pm 2.33~\mathrm{km~s^{-1}~Mpc^{-1}},~r_{\mathrm{d}}=144.7^{+4.5}_{-5.2}~\mathrm{Mpc}$. The constrains of $p_{\mathrm{age}},~H_{0},~M_{\mathrm{B}},~r_{\mathrm{d}}$ agree with PAge, except for $\eta~(\eta_{1})$, which shifts toward more negative values and has a larger uncertainty in MAPAge, caused by the added cubic term. Furthermore, the reconstructed BAO distances $D_{\mathrm{H}}/r_{\mathrm{d}},~D_{\mathrm{M}}/r_{\mathrm{d}},~D_{\mathrm{v}}/r_{\mathrm{d}}$ are shown in Fig. 5. The results exhibit no significant difference from PAge in either central values or uncertainties. The enlarged uncertainty in $\eta~(\eta_{1})$ minimally affects $D_{\mathrm{H}}/r_{\mathrm{d}},~D_{\mathrm{M}}/r_{\mathrm{d}},~D_{\mathrm{v}}/r_{\mathrm{d}}$, demonstrating stability with respect to the chosen truncated order in Eq. (1).
To quantitatively evaluate the performance of the PAge model relative to the $\Lambda\mathrm{CDM}$ model, we perform Bayesian inference across different datasets. The MAPAge does not show significant different with PAge, as demonstrated by Bayesian inference. Therefore, for simplicity , we just focus on PAge model. Specifically, the Bayes factors $\mathrm{ln}\mathcal{B}_{ij}$ are calculated in logarithmic space,
@@
where $P(D\mid\theta,M)$ is the likelihood of the data $D$ given the parameters $\theta$ and model $M$, $P(\theta\mid M)$ is the prior probability of $\theta$ under $M$, and $P(M)$ is the prior of $M$. The extent of model preference is assessed using the Jeffreys scale, as described in Ref. Trotta (2008).
@@
  • $1\leq|\mathrm{ln}\mathcal{B}_{ij}|<2.5$, weak evidence.
  • $2.5\leq|\mathrm{ln}\mathcal{B}_{ij}|<5$, moderate evidence.
  • $5\leq|\mathrm{ln}\mathcal{B}_{ij}|<10$, strong evidence.
@@
Given the impact of prior width on our results, we follow Refs. Patel et al. (2024); Sola Peracaula et al. (2019) and broaden the original prior width $d$ by a factor $\alpha$, so that the new prior width is $d^{*}=\alpha d$. In the case of uniform priors Patel et al. (2024); Sola Peracaula et al. (2019), the new Bayes factors $\ln\mathcal{B}^{*}_{ij}$ are modified into
$\ln\mathcal{B}_{ij}^{*}=\ln\left[\frac{(\alpha_{\nu_{1}}\alpha_{\nu_{2}}\alpha_{\nu_{3}}\dots)_{j}}{(\alpha_{\nu_{1}}\alpha_{\nu_{2}}\alpha_{\nu_{3}}\dots)_{i}}\right]+\ln\mathcal{B}_{ij},$ (28)
where $(\alpha_{\nu_{1}}\alpha_{\nu_{2}}\alpha_{\nu_{3}}\dots)_{i}$ denotes the product of prior-width broadening factors $\alpha_{\nu_{k}}$ for parameters $\nu_{k}$ in PAge and $(\alpha_{\nu_{1}}\alpha_{\nu_{2}}\alpha_{\nu_{3}}\dots)_{j}$ denotes the corresponding product for $\Lambda\mathrm{CDM}$. We consider only the impact of the prior widths of $\eta,~p_{\mathrm{age}}$ in PAge. For simplicity, we set $\alpha_{\eta}=~\alpha_{p_{\mathrm{age}}}=\alpha$. So the Eq. (28) becomes
$\ln\mathcal{B}_{ij}^{*}=\ln\mathcal{B}_{ij}-2\ln\alpha.$ (29)
In the left panel of Fig. 6, we present the final model comparison results, with the PAge model assessed relative to the fiducial $\Lambda\mathrm{CDM}$ across four datasets. Positive values indicate a preference for the PAge model, while negative values indicate a preference for the $\Lambda\mathrm{CDM}$ model. Specifically, when $\alpha=1$ we find $\mathrm{ln}\mathcal{B}_{ij}=9.17,~4.81,~1.01,~\text{and}~2.71$ for DESY5+DESI DR2+CC, DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC, respectively. It is clear that the PAge model consistently provides a better fit than the standard $\Lambda\mathrm{CDM}$ across all datasets. In particular, for the DESY5+DESI DR2+CC datasets, the Bayes factor is 9.17, indicating nearly decisive evidence in favor of the PAge model over the $\Lambda\mathrm{CDM}$ model. Varying $\alpha$ can switch the preferred model. When $\alpha\geq 12,~2,~4$ for DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC, respectively, the Bayes factor favors $\Lambda\mathrm{CDM}$. By contrast, for DESY5+DESI DR2+CC, the preference flips for $\alpha\geq 98$.
Moreover, to assess the performance of PAge compared to $w_{0}w_{a}\mathrm{CDM}$, we perform Bayesian inference in the same manner. The prior ranges used for $w_{0}w_{a}\mathrm{CDM}$ in this work are $w_{0}\in[-3,1],~w_{a}\in[-3,2],~M_{\mathrm{B}}\in[-21,-19],~H_{0}\in[60,80]~\mathrm{km~s^{-1}~Mpc^{-1}},~r_{\mathrm{d}}\in[120,180]~\mathrm{Mpc},~\Omega_{\mathrm{m}}\in[0.2,0.4]$. We also explore the impact of prior width by broadening the prior widths on $\eta$ and $p_{\mathrm{age}}$. The Bayes factors take the same form as Eq. (29), where $i$ denotes PAge and $j$ denotes $w_{0}w_{a}\mathrm{CDM}$. The final results are shown in the right panel of Fig. 6. For $\alpha=1$, we find $\ln\mathcal{B}_{ij}=-1.25,~0.32,~1.31,~\text{and}~2.41$ for DESY5+DESI DR2+CC, DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC, respectively. These results indicate at most weak evidence in favor of $w_{0}w_{a}\mathrm{CDM}$ model over PAge model. For DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC, the evidence shifts in favor of $w_{0}w_{a}\mathrm{CDM}$ when $\alpha\geq 2,~2,~4$, respectively. By contrast, for DESY5+DESI DR2+CC, PAge is favored only when $\alpha\leq 0.5$. These conclusions also depend on the chosen prior width for $w_{0},~w_{a}$, which are held fixed here for simplicity. Widening those prior widths on $w_{0}$ and $w_{a}$ tends to strengthen the evidence in favor of the PAge model.
@@ IV CONCLUSION
The tension between the early-Universe and late-Universe measurements of $H_{0}$ has become one of the most significant challenges in modern cosmology. Recently, DESI DR1 Adame et al. (2025) and DR2 Abdul Karim et al. (2025) have further sharpened this tension with addition hints of dynamical dark energy in $3.9\sigma$ and $4.2\sigma$ level. In the context of the $H_{0}$ tension, numerous theoretical models have been proposed, although many models face corresponding no-go arguments. Similarly, the hint of dynamical dark energy has been critically examined through the reanalyzes of the DESY5 SNe data Huang et al. (2025b); Efstathiou (2025); Vincenzi et al. (2025); Gialamas et al. (2025b) and the DESI BAO measurements Zheng et al. (2025); Mukherjee and Sen (2025); Liu et al. (2024).
To provide a model-independent perspective, we employ the PAge-improved IDL in this work to constrain cosmological parameters. Our work is based on four datasets, namely DESY5+DESI DR2+CC, DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC, allowing us to investigate the respective roles of SNe and BAO data in constraining cosmological parameters. For DESY5+DESI DR2+CC, we find $H_{0}=67.91\pm 2.33~\mathrm{km~s^{-1}~Mpc^{-1}}$, which agrees with the Planck 2018 result, but exhibits a mild $2.0\sigma$ tension with the SH0ES measurement. The other datasets also remain consistent with Planck 2018. Compared to SH0ES, the resulting tensions are $2.1\sigma$, $1.6\sigma,$ and $1.7\sigma$ for DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC, respectively. In our results, the DESY5 data tend to favor a lower value of $H_{0}$ and a higher value of $M_{\mathrm{B}}$, which correspond to a shift in the intercept of the magnitude–redshift relation within the $H_{0}$$M_{\mathrm{B}}$ parameter space. Additionally, the DESI DR2 data indicate a preference for a higher value of $r_{\mathrm{d}}$ compared to SDSS data, with a difference of approximately 2 Mpc.
@@
Mapping specific cosmological models into the PAge parameter space $(p_{\mathrm{age}},\eta)$, our results align with the DESI DR2 hints of dynamical dark energy and deviate from $\Lambda\mathrm{CDM}$. In particular, the reconstructed BAO distance of the PAge model reveals a special trend compared to best-fit $\Lambda\mathrm{CDM}$. It shows a slight enhancement in very low redshift, suppression at intermediate redshift, and an upturn at high redshift, which may suggest that the $\Lambda\mathrm{CDM}$ model is insufficient to fully reconcile the observed data. Furthermore, the Bayesian model comparison provides evidence in favor of the PAge model over $\Lambda\mathrm{CDM}$, with Bayes factors of 9.17, 4.81, 1.01, and 2.71 in DESY5+DESI DR2+CC, DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, PantheonPlus+SDSS+CC respectively. It gives nearly decisive evidence in favor of the PAge model over $\Lambda\mathrm{CDM}$ in the DESY5+DESI DR2+CC combination. Comparing PAge with $w_{0}w_{a}\mathrm{CDM}$, the Bayes factors are 1.25,0.32,1.31,-1.25,~0.32,~1.31, and $2.41$ for DESY5+DESI DR2+CC, DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC, respectively, showing at most weak evidence in favor of $w_{0}w_{a}\mathrm{CDM}$. The interpretation of these comparisons is influenced by the adopted prior width. The wider prior width for PAge parameters, the weaker the evidence in favor of PAge.
@@
In future work, we aim to explore the data in more detail by performing a redshift-binned analysis. Motivated by the connection between the expansion history and the dark-energy equation of state, we will further investigate the occurrence of unphysical phantom behaviors Caldwell et al. (2003); Vikman (2005) using the PAge model. It should be noted that the PAge model is essentially a Taylor expansion, which may suffer from singularities as discussed in Ref. Zhang et al. (2017b) when reconstructing the equation-of-state parameter. We are going to calibrate the Taylor expansion with alternatives beyond the early-matter-dominated Universe used in the PAge model, to avoid the singularities. These efforts may provide new insights into the nature of dark energy and further clarify the physical origin of the $H_{0}$ tension.